OpenStudy (anonymous):

The sum of three consecutive integers is 27. Find the integers.

OpenStudy (anonymous):

8 9 10

OpenStudy (anonymous):

x+(x+1)+(x+2)=27 Remove the parentheses that are not needed from the expression. x+x+1+x+2=27 Since x and x are like terms, add x to x to get 2x. 2x+1+x+2=27 Since 2x and x are like terms, add x to 2x to get 3x. 3x+1+2=27 Add 2 to 1 to get 3. 3x+3=27 Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides. 3x=-3+27 Add 27 to -3 to get 24. 3x=24 Divide each term in the equation by 3. (3x)/(3)=(24)/(3) Simplify the left-hand side of the equation by canceling the common terms. x=(24)/(3) Simplify the right-hand side of the equation by simplifying each term. x=8 x+1 = 9 x+2 = 10

OpenStudy (anonymous):

i need a solution on this...

OpenStudy (anonymous):

i gave u leinyr 0909

OpenStudy (anonymous):

ahh i already read it..sorry haha

OpenStudy (anonymous):

haha , no problem

OpenStudy (anonymous):

checking please..thanks

OpenStudy (anonymous):

i need a checking for this please answer